OpenStudy (anonymous):

1. The inequality x ≥ -3 is satisfied by point (0, -2). True False. 2.The inequality 5 x - 2 y + 1 > 0 is satisfied by point (-2, -1). True False. 3. The inequality 2 x - 3 y ≥ 5 is satisfied by point ( 1 over 2 , 1 over 3 ). True False

OpenStudy (texaschic101):

x >= -3 .......(0,-2) 0 >= -3 TRUE 5x - 2y + 1 > 0.......(-2,-1) 5(-2) - 2(-1) + 1 > 0 -10 + 2 + 1 > 0 -7 > 0 FALSE 2x - 3y >= 5 .....(1/2, 1/3) 2(1/2) - 3(1/3) >= 5 1 - 1 >= 5 0 >= 5 FALSE

OpenStudy (anonymous):

Thank you!

OpenStudy (texaschic101):

anytime :)